# Condition Weaving Meets Expert Modulation: Towards Universal and Controllable Image Generation

### v1

@@ Abstract
1 Introduction
Figure 1: The UniGen architecture supports image inputs with arbitrary types of conditioning to control the content of image generation.
@@ 1 Introduction
With the continuous advancement of diffusion-based methods [1, 2, 3, 4, 5], model architectures have evolved from the early UNet [6, 1] design to the more recent DiT [7] structure, leading to significant improvements in the quality of synthesized images, particularly in text-guided image generation tasks. However, relying solely on text-to-image generation often fails to meet the requirements for controllable synthesis in scenarios demanding precise spatial or structural guidance. To address this limitation, image-to-image generation based on conditional visual constraints has emerged as another important branch of image synthesis. This task aims to generate images under spatial constraints using various conditional inputs, such as depth, canny, and pose information, etc.
@@
To tackle this task, ControlNet [8] introduces a separate branch to inject spatial condition information into the generation process. IP-Adapter [9] incorporates an adapter structure to enable partial parameter tuning while enforcing spatial constraints. Subsequently, several works [10, 11] have explored approaches inspired by mixture-of-experts (MoE) architectures to handle diverse types of visual conditions. More recently, with the rapid development of the FLUX [5] model in image synthesis, a number of studies [12, 13, 14] have adopted lightweight fine-tuning techniques such as LoRA [15] to enhance spatial constraint capabilities from conditional images by tuning only a subset of model parameters.
@@
However, existing methods—whether through constructing MoE architectures or introducing additional parameter tuning—typically design independent control units or control branches for each conditional visual input. Therefore, two major limitations remain in diverse conditional image generation tasks: Firstly, redundant parameters and inefficient computation. On one hand, some approaches build independent and parallel control units for each condition, leading to significant parameter redundancy. On the other hand, certain methods adopt MoE-like designs that merely select specific preprocessing modules based on the condition type, which substantially differs from mixture-of-experts architectures designed for efficient parallel computation. These strategies overlook both the significant differences and the potential commonalities among conditions, heavily relying on condition-specific modules. This leads to excessive parameter redundancy and inefficient computation, ultimately limiting the model’s practicality. Secondly, a lack of effective interaction between the global visual representations guided by text and the local visual representations guided by conditional images. Existing methods typically rely on isolated conditional control modules to extract conditional visual representations, failing to exploit the dynamic interplay between prompt-guided global (holistic) features and condition-image-guided local (spatial) features. This lack of interaction introduces semantic discrepancies between global and local visual cues, often leading to visually abrupt or inconsistent synthesis results.
@@
To address the above challenges, we propose a unified image generation framework, as illustrated in Fig. 1. Firstly, to address the significant computational redundancy caused by existing methods that employ separate modules or branches for processing different conditional visual representations, we propose the Condition Modulated Expert (CoMoE) module. This method starts by analyzing the differences and commonalities among various types of conditional visual representations. It aggregates patch features from conditional visual inputs into foreground and background regions based on feature similarity, then routes them to dedicated Modulated Expert (MoE) modules for feature filtering and aggregation according to the condition type. This design eliminates the need for separate preprocessing modules for each condition type, thereby reducing parameter redundancy. Furthermore, to bridge the semantic gap between global visual representations of the base model and local spatial representations from the conditional control units, caused by their independent processing, we propose the WeaveNet architecture. By employing a streamlined serpentine data flow, WeaveNet effectively integrates prompt-guided global semantic constraints with condition image-guided local spatial constraints, enhancing their interaction and significantly mitigating discrepancies arising from separately processing prompts and conditional images. To validate the effectiveness of our approach, we conduct experiments on the MultiGen-20M [16] and Subjects-200K [13] datasets. Compared to both traditional and latest methods, our approach achieves superior performance across multiple evaluation metrics, with only a slight performance gap in certain metrics compared to methods built upon the FLUX [5] base model.
@@
2 Related Works
@@
Controllable image generation aims to guide image synthesis by jointly leveraging conditional images and textual prompts. Early methods (e.g., [8]) align spatial features by injecting condition information via parallel branches into the main network. Unicontrol [10] employs a Mixture-of-Experts (MoE)-like architecture [17, 18] that selects pretrained embedding layers based on the condition type to achieve feature alignment. Recent works leverage FLUX [5] as a strong backbone, fine-tuning it with LoRA [15] modules for different condition types (e.g., [13, 12, 14]). Despite their effectiveness, these approaches often require training separate LoRA or ControlNet modules or adding parallel branches, which increases model size and complexity, and limits generalization. Moreover, handling each condition independently introduces redundant parameters and overlooks potential shared representations across different conditions. To address these challenges, this paper analyzes the commonalities and distinctions among visual representations of different condition types and proposes an approach that aggregates similar features using independent modules to reduce redundant computation. Additionally, an interactive injection strategy is introduced to align globally text-guided visual features with locally condition-guided features, thereby bridging the semantic gap between them.
3 Method
Figure 2: The architecture of Condition Modulated Experts (CoMoE) Module.
@@ 3 Method
To better support visual control under diverse conditional inputs and enable a unified image-to-image generation framework, we propose UniGen. As illustrated in Fig. 1, UniGen accommodates image generation guided by any of 12 types of conditional inputs. During training, it follows the computational framework of existing image-to-image generation methods [13, 10] and performs conditional noise prediction using a Flow Matching-based [19] approach. First, to address the computational and parameter redundancy caused by existing methods that employ parallel and independent modules for different condition types, we introduce the Condition Modulation Expert (CoMoE) module. As shown in Fig. 2, we leverage the commonalities and differences across diverse condition types to guide expert module selection based on the similarity between patch representations. For patches within the same condition type, we independently process the similar representations of foreground regions, aiming to minimize parameter redundancy. Second, to tackle the lack of effective interaction between global visual representations guided by text and local spatial visual representations guided by conditional images, we introduce the WeaveNet architecture. As illustrated in Fig. 3(b), WeaveNet enhances interactive control between text and conditional images through a staggered serpentine computation flow, mitigating local semantic discrepancies and inconsistencies.
3.1 Condition Modulated Experts (CoMoE)
@@
To fully exploit the diversity and consistency among various types of conditional visual representations while reducing computational and parameter redundancy, we propose a Conditional Modulation Expert (CoMoE) module. As illustrated in Fig. 2, inspired by the design of Mixture of Experts (MoE) commonly used in large multimodal models [20, 21], the module is composed of independent experts and shared experts. The independent experts are designed to capture common foreground information within the same type of conditional representations, enabling joint representation learning of conditional semantics and visual features for condition-specific visual preprocessing. The shared experts focus on integrating global visual representations guided by textual prompts with local conditional visual features, ensuring consistency between condition information and prompt guidance while reducing their semantic gap.
@@
First, to address representation ambiguity, information loss, and misalignment caused by certain sparsely structured visual conditions (e.g., Canny, HED, HEDSketch), we introduce a text-guided global visual feature $F_{n}$. By fusing the global representation $F_{n}$ with the condition-specific representation $F_{c}$, we jointly determine the selection of independent expert modules, allowing better control over the denoising direction, as shown in Equation 1. Both $F_{n}$ and $F_{c}$ are preprocessed by a VAE [22] and then converted into token-level visual representations using patch embedding. During training, noise is added to the encoded visual representations $F_{n}$ with a strength factor $\sigma$ as follows: $F_{n}=\sigma\cdot N+(1-\sigma)\cdot F_{n}$, where $N$ denotes the noise. During inference, $F_{n}=N$ is used directly as the input to the diffusion model.
@@
Secondly, to address the loss and misalignment of spatial information caused by Modulated Expert modules processing patch features at the token level, we incorporate Rotary Positional Encoding (RoPE) [23] to embed spatial cues into visual features. In the Modulated Expert module, we aim to leverage condition-specific semantics to process localized visual representations and fuse them with spatially aligned, text-guided global visual features. Specifically, we construct a learnable parameter matrix $LW_{c}$, and extract the semantic representation $E_{c}$ based on the category of the given condition. The condition-aware semantic embedding is then used to modulate the parameter matrix, enabling one distinct matrix per condition for conditional feature representation, as shown in Equation 2.
@@
In the above equation, $\cdot$$\cdot$] denotes the token-level concatenation of features. $F_{p}$ denotes the semantic representation of the prompt encoded by CLIP [24] and T5 [25].
@@
3.2 WeaveNet
(a) The architecture of ControlNet.
(b) The architecture of WeaveNet.
Figure 3: (a) illustrates the structure of the traditional ControlNet, while (b) shows our proposed WeaveNet architecture, which performs conditional control in an interwoven manner. The Zero ResProjection module, initialized with zeros, is a residual network designed to fuse the noise-injected visual representation with the conditional representation.
To address the semantic gap between textual prompts and conditional visual control introduced by existing methods such as ControlNet [8], which processes conditional visual features and textual guidance in parallel via independent branches, we propose a novel architecture named WeaveNet. As illustrated in Fig. 3(b), WeaveNet introduces an interactive conditional information injection mechanism that enables dynamic interaction between conditional visual representations and text-guided visual features. Specifically, following the architectural design of ControlNet, we replicate the modules and parameters of the base model to construct a separate branch for processing conditional inputs. Unlike ControlNet, WeaveNet integrates features from the base model into the corresponding layers of the control branch, effectively combining the global semantics from textual prompts with the local information from conditional images. Then, the Zero ResProjection Module is employed to align the outputs of each layer in the control branch and inject them into the main network, ensuring dynamic interaction between the prompt and the conditional image. This helps prevent semantic gaps caused by overemphasis on either side. The detailed computational process is presented in Algorithm 1.
@@ 3.2 WeaveNet
4 Experiments
4.1 Datasets
We conduct training and evaluation on the MultiGen-20M [16] and Subjects-200K [13]datasets. MultiGen-20M contains over 2 million image-description pairs with corresponding conditional images. These conditional images cover 12 types of conditions, such as depth, Canny edges, OpenPose, HED, normals, and grayscale. However, due to significant inconsistencies in the descriptions constructed for each image in the dataset and generally low annotation accuracy, we re-annotated the entire dataset using Qwen [26] for model training and evaluation. Subjects-200K consists of over 200,000 high-quality images paired with detailed entity annotations and accurate textual descriptions, offering higher image fidelity and more precise captions compared to MultiGen-20M. To further enrich Subjects-200K, we generate conditional images using Dwpose [27] (for OpenPose), Depth Anything [28] (for depth), and Canny [29] operator (for edge maps). For both datasets, we randomly sample 5,000 image-description-condition triplets to construct the test set. We will upload the newly annotated data and test set information to Hugging Face in the future.
4.2 Evaluation Metrics

### v3

@@ Abstract
I Introduction
Figure 1: (a) Overall architecture of UniGen, which employs Modulated Experts for conditional modulation and feature alignment to reduce redundant computation, and integrates textual and visual control signals through the interactive WeaveNet design. (b) MoE-based architectures used in related methods [Unicontrol], where separate preprocessing units are designed for each condition. (c) Performance comparison showing UniGen’s improvements over UniControl [Unicontrol] in parameters, inference speed, and evaluation metrics.
@@ I Introduction
With the continuous advancement of diffusion-based methods [ddpm, improved_ddpm, stable_diffusion, sd35, flux.1], model architectures have evolved from the early UNet [unet, ddpm] design to the more recent DiT [dit] structure, leading to significant improvements in the quality of synthesized images, particularly in text-guided image generation tasks. However, relying solely on text-to-image generation often fails to meet the requirements for controllable synthesis in scenarios demanding precise spatial or structural guidance. To address this limitation, image-to-image generation based on conditional visual constraints has emerged as another important branch of image synthesis. This task aims to generate images under spatial constraints using various conditional inputs, such as depth, canny, and pose information, etc.
@@
To tackle this task, ControlNet [controlnet] introduces a separate branch to inject spatial condition information into the generation process. IP-Adapter [ipadapter] incorporates an adapter structure to enable partial parameter tuning while enforcing spatial constraints. Subsequently, several works [Unicontrol, Pixelponder] have explored approaches inspired by mixture-of-experts (MoE) architectures to handle diverse types of visual conditions. More recently, with the rapid development of the FLUX [flux.1] model in image synthesis, a number of studies [Unicombine, ominicn, Easycontrol] have adopted lightweight fine-tuning techniques such as LoRA [lora] to enhance spatial constraint capabilities from conditional images by tuning only a subset of model parameters.
@@
However, existing methods—whether through constructing MoE architectures or introducing additional parameter tuning—typically design independent control units or control branches for each conditional visual input. Therefore, two major limitations remain in diverse conditional image generation tasks: Firstly, redundant parameters, inefficient computation and information forgetting under multi-condition control. As shown in Fig. 1 (b), some methods [Unicontrol] adopt an MoE-like design [moe_1] that selects specific preprocessing modules solely based on the input condition type. This differs substantially from mixture-of-experts architectures, which are designed for efficient parallel computation. In addition, certain approaches [ominicn] build independent and parallel control units for each condition, resulting in severe parameter redundancy. These methods share a fundamental limitation: they overlook the substantial differences between various conditions as well as the underlying commonalities among them. These approaches rely heavily on condition-specific designs, where each condition type is assigned an independent processing module for feature integration. Such designs introduce substantial parameter redundancy and greatly reduce computational efficiency, thereby constraining the model’s practical utility. Moreover, using isolated modules for each condition type prevents effective alignment of information when handling tasks involving multiple conditions. The model struggles to aggregate heterogeneous conditional cues and suffers from severe information-forgetting problems, making it difficult to establish meaningful correlations among multiple image-level conditions. As a result, the model fails to achieve coherent joint control across conditions, which significantly limits its scalability and generalization ability. Secondly, a lack of effective interaction between the global visual representations guided by text and the local visual representations guided by conditional images. Existing methods typically rely on isolated conditional control modules to extract conditional visual representations, failing to exploit the dynamic interplay between prompt-guided global (holistic) features and condition-image-guided local (spatial) features. This lack of interaction introduces semantic discrepancies between global and local visual cues, often leading to visually abrupt or inconsistent synthesis results.
@@
To address the above challenges, we propose a unified image generation framework, as illustrated in Fig. 1 (a). Firstly, to address the significant computational redundancy caused by existing methods that employ separate modules or branches for processing different conditional visual representations, we propose the Condition Modulated Expert (CoMoE) module. Our method abandons the traditional paradigm in which independent, non-interfering modules are designed for different condition types. Instead, we start from the common nature of conditional visual representations and focus on the shared foreground-related information across visual tokens. By uncovering and aggregating the latent commonalities among different visual tokens, we employ a conditional modulated expert to fuse token-level visual features with semantic conditions. Compared with traditional approaches, our method not only supports high-quality image synthesis under single-condition control, but also extends easily and effectively to multi-condition joint control, enabling high-quality image generation under multiple conditioning inputs. Moreover, as shown in Fig. 1(c), compared with UniControl [Unicontrol], which adopts an MoE-like architecture, our method significantly improves image quality while reducing both parameter count and inference latency by approximately 50%. At the same time, it enables joint control under multiple input conditions. Furthermore, to bridge the semantic gap between global visual representations of the base model and local spatial representations from the conditional control units, caused by their independent processing, we propose the WeaveNet architecture. By employing a streamlined serpentine data flow, WeaveNet effectively integrates prompt-guided global semantic constraints with condition image-guided local spatial constraints, enhancing their interaction and significantly mitigating discrepancies arising from separately processing prompts and conditional images. To validate the effectiveness of our approach, we conduct experiments on the MultiGen-20M [cn++] and Subjects-200K [ominicn] datasets. Compared to both traditional and latest methods, our approach achieves superior performance across multiple evaluation metrics, with only a slight performance gap in certain metrics compared to methods built upon the FLUX [flux.1] base model.
@@
II Related Works
II-A Diffusion Models
With the rapid development of diffusion models in the field of image generation, they have gradually emerged as a new generative modeling paradigm following GANs. Initially proposed by [diffusion_model], diffusion models add noise to original data through a forward process and learn the reverse process to reconstruct the original image. Denoising Diffusion Probabilistic Models (DDPM) [ddpm] significantly improved sampling stability and generation quality. To further accelerate the sampling process and enhance output quality, [ddim] introduced a non-Markovian diffusion process. [improved_ddpm] incorporated multiple training strategies to boost model performance. SDE-based diffusion models [SDE] unified the diffusion framework from the perspective of stochastic differential equations and proposed gradient-based sampling strategies, enhancing model flexibility. In recent years, with the widespread adoption of Transformer architectures in diffusion models, models such as DiT [dit] have replaced the traditional U-Net [unet] backbone and achieved remarkable progress in high-resolution image generation tasks. Furthermore, the emergence of the Stable Diffusion [stable_diffusion, sd35] and FLUX [flux.1] models has led to substantial improvements in text-controlled image generation. However, considering that the FLUX series involves larger parameter sizes and higher complexity than the Stable Diffusion models, we adopt Stable Diffusion as the backbone of our proposed UniGen architecture to achieve controllable image generation with greater efficiency.
II-B Controllable Image Generation
Controllable image generation aims to guide image synthesis by jointly leveraging conditional images and textual prompts. Existing approaches can be broadly categorized into two groups. The first category focuses on generating semantically aligned content with the text prompt, without requiring strict spatial alignment with the conditional image. For example, [ipadapter] introduces an adapter mechanism that aggregates visual and textual inputs to jointly constrain the generation process. [blip_diff] leverages a multimodal vision-language model to fuse conditional images and semantic information, which are then fed into a diffusion model for controllable generation. [dreambooth, dreamrender] employs a Multi-Modal Attention (MM Attention) mechanism to enable targeted condition injection and personalized image synthesis. The second category emphasizes preserving both semantic and spatial consistency with the conditional image. Early works such as [controlnet] achieve spatial alignment by encoding conditional information through a parallel branch and integrating it with the text-guided visual features from the backbone network. [Unicontrol] extends this idea by designing a Mixture-of-Experts (MoE)-like architecture [moe_1, moe_2] that supports controllable generation from diverse input conditions. More recently, FLUX [flux.1] has emerged as a strong backbone for text-controlled image generation. Several works [ominicn, Unicombine, Easycontrol] fine-tune FLUX with LoRA [lora] modules on different conditional image types to achieve controllable generation. [Pixelponder] takes a different approach by introducing an MoE Adapter and Patch Adaption module to align multi-condition visual features at the preprocessing stage.
@@ II-B Controllable Image Generation
Although these methods demonstrate strong controllability, they often rely on training multiple LoRA [lora] or ControlNet [controlnet] structures or designing parallel preprocessing modules for visual condition alignment. This increases both the number of control parameters and model complexity, significantly limiting generalization. Moreover, independently handling each condition leads to redundant parameters and ignores shared representations among conditional inputs. To address these challenges, this paper analyzes the commonalities and distinctions among visual representations of different condition types and proposes an approach that aggregates similar features using independent modules to reduce redundant computation. Additionally, an interactive injection strategy is introduced to align globally text-guided visual features with locally condition-guided features, thereby bridging the semantic gap between them.
III Method
(a) The UniGen architecture supports image inputs with arbitrary types of conditioning to control the content of image generation.
(b) The architecture of Condition Modulated Experts (CoMoE) Module.
Figure 2: (a) Overall architecture of UniGen, enabling controllable image generation with arbitrary conditional inputs. (b) Structure of Condition Modulated Experts (CoMoE), which modulates and aligns features based on conditional information, achieving parameter-efficient and high-quality controllable generation through a unified branch.
@@ III Method
To better support visual control under diverse conditional inputs and enable a unified image-to-image generation framework, we propose UniGen. As illustrated in Fig. 2(a), UniGen accommodates image generation guided by any of 12 types of conditional inputs. During training, it follows the computational framework of existing image-to-image generation methods [ominicn, Unicontrol] and performs conditional noise prediction using a Flow Matching-based [flow_matching] approach. First, to address the computational and parameter redundancy caused by existing methods that employ parallel and independent modules for different condition types, we introduce the Condition Modulation Expert (CoMoE) module. As shown in Fig. 2(b), we leverage the commonalities and differences across diverse condition types to guide expert module selection based on the similarity between patch representations. For patches within the same condition type, we independently process the similar representations of foreground regions, aiming to minimize parameter redundancy. Second, to tackle the lack of effective interaction between global visual representations guided by text and local spatial visual representations guided by conditional images, we introduce the WeaveNet architecture. As illustrated in Fig. 3, WeaveNet enhances interactive control between text and conditional images through a staggered serpentine computation flow, mitigating local semantic discrepancies and inconsistencies.
III-A Condition Modulated Experts (CoMoE)
@@
To fully exploit the diversity and consistency among various types of conditional visual representations while reducing computational and parameter redundancy, we propose a Conditional Modulation Expert (CoMoE) module. As illustrated in Fig. 2(b), inspired by the design of Mixture of Experts (MoE) commonly used in large multimodal models [moe_llava, remoe], the module is composed of independent experts and shared experts. The independent experts are designed to capture common foreground information within the same type of conditional representations, enabling joint representation learning of conditional semantics and visual features for condition-specific visual preprocessing. The shared experts focus on integrating global visual representations guided by textual prompts with local conditional visual features, ensuring consistency between condition information and prompt guidance while reducing their semantic gap.
@@
First, to address representation ambiguity, information loss, and misalignment caused by certain sparsely structured visual conditions (e.g., Canny, HED, HEDSketch), we introduce a text-guided global visual feature $F_{n}$. Note that $F_{n}$ denotes the features after the first Transformer Block of the backbone network(DiT), where an initial stage of vision–text fusion has already been performed. By fusing the global representation $F_{n}$ with the condition-specific representation $F_{c}$, we jointly determine the selection of independent expert modules, allowing better control over the denoising direction, as shown in Equation 1. Both $F_{n}$ and $F_{c}$ are preprocessed by a VAE [vae] and then converted into token-level visual representations using patch embedding. During training, noise is added to the encoded visual representations $F_{n}$ with a strength factor $\sigma$ as follows: $F_{n}=\sigma\cdot N+(1-\sigma)\cdot F_{n}$, where $N$ denotes the noise. During inference, $F_{n}=N$ is used directly as the input to the diffusion model.
@@
Secondly, when performing subsequent independent processing of vision-level tokens, all visual tokens are first dispersed and then passed to separate Modulated Expert modules based on their feature correlations. This procedure, however, may cause the loss of global structural and spatial information from the original conditioning image. Therefore, before discretizing the tokens, we introduce Rotary Positional Embeddings (RoPE) [rope] to encode spatial positional information, ensuring spatial consistency and preventing spatial confusion in the aggregated visual features after independent expert processing (we provide experimental analysis of RoPE in Section IV-F2). In addition, within the Modulated Expert module, our goal is to leverage condition-specific semantics to process local visual representations and fuse them with spatially aligned, text-guided global visual features. Specifically, we construct a learnable parameter matrix $LW_{c}$, and extract the semantic representation $E_{c}$ based on the category of the given condition. The condition-aware semantic embedding is then used to modulate the parameter matrix, enabling one distinct matrix per condition for conditional feature representation, as shown in Equation 2.
@@
Figure 3: WeaveNet architecture diagram: conditional control information is injected by interleaving the backbone network with conditional-control units.
@@
Although this design aims to aggregate homogeneous conditional visual information and perform directional, condition-guided feature enhancement to achieve discretized feature alignment, it inevitably suffers from a common issue in expert-based architectures: a few Modulated Expert modules may be overly activated, dominating the modulation of most visual tokens. This results in dense computation despite the intended sparsity. Meanwhile, our Modulated Expert modules are designed to fuse and process visually similar tokens based on conditional semantic representations. Excessive activation of a single Modulated Expert increases training difficulty and harms convergence. To address this, we introduce an expert regularization loss [moe_loss] widely used in large language models. By minimizing the inner product between expert scores and expert activation counts, the loss prevents the router from over-assigning tokens to a small subset of experts, thereby encouraging balanced expert utilization. In addition, this regularization effectively prevents visual tokens from being overly routed to a single Modulated Expert, ensuring that each expert focuses on a more coherent subset of visual representations and accelerating convergence during training.
@@
In the above equation, $\cdot$$\cdot$] denotes the token-level concatenation of features. $F_{p}$ denotes the semantic representation of the prompt encoded by CLIP [clip] and T5 [t5].
@@
III-B WeaveNet
To address the semantic gap between textual prompts and conditional visual control introduced by existing methods such as ControlNet [controlnet], which processes conditional visual features and textual guidance in parallel via independent branches, we propose a novel architecture named WeaveNet. As illustrated in Fig. 3, WeaveNet introduces an interactive conditional information injection mechanism that enables dynamic interaction between conditional visual representations and text-guided visual features. Specifically, following the architectural design of ControlNet, we replicate the modules and parameters of the base model to construct a separate branch for processing conditional inputs. Unlike ControlNet, WeaveNet integrates features from the base model into the corresponding layers of the control branch, effectively combining the global semantics from textual prompts with the local information from conditional images. Then, the Zero ResProjection Module is employed to align the outputs of each layer in the control branch and inject them into the main network, ensuring dynamic interaction between the prompt and the conditional image. This helps prevent semantic gaps caused by overemphasis on either side. The detailed computational process is presented in Algorithm 1.
@@ III-B WeaveNet
IV Experiments
IV-A Datasets
We conduct training and evaluation on the MultiGen-20M [cn++] and Subjects-200k [ominicn]datasets. MultiGen-20M contains over 2 million image-description pairs with corresponding conditional images. These conditional images cover 12 types of conditions, such as depth, Canny edges, OpenPose, HED, normals, and grayscale. However, the image descriptions in this dataset exhibit significant inconsistencies and low accuracy, which introduces training bias and degrades generation quality. To address this issue, we re-annotated all captions in the dataset using Qwen [^1] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct [qwen]. The newly annotated captions have been uploaded to Hugging Face [^2] https://huggingface.co/datasets/gavin-zhang/MultiGen20M_json. Furthermore, to ensure fair and accurate comparisons, all methods are trained and evaluated using the same training and test splits. Compared with MultiGen-20M, the Subjects-200k dataset contains over 200k high-quality images with detailed and accurate entity-level annotations and reliable textual descriptions, along with GPT-based scoring of the captions. However, Subjects-200k provides relatively weaker support for conditional image configurations. To overcome this limitation, we further extend Subjects-200k by generating three types of conditional inputs: Depth, Canny, and OpenPose. Specifically, we use the pre-trained Depth-Anything [^3] https://github.com/LiheYoung/Depth-Anything [depthanything] model to produce depth maps for each image, and the pre-trained DWpose [^4] https://github.com/IDEA-Research/DWPose [dwpose] model to extract human keypoints (OpenPose). For the Canny [canny] condition, we generate edge maps online using the OpenCV-Python library. Note that DWpose may incorrectly produce pose information for images without humans. To avoid this, we apply an additional filtering step using Qwen: for each image, we determine whether it contains human subjects, and remove the corresponding OpenPose condition if no humans are present. All extended data have been uploaded to Hugging Face [^5] https://huggingface.co/datasets/gavin-zhang/Subjects200K. We also strictly separate the training and test sets and retrain all compared methods to ensure fair and valid evaluation. For both MultiGen-20M and Subjects-200k, the test set consists of 5,000 condition–description–reference image triplets, kept entirely independent from the training data. In addition, since Subjects-200k supports both single-condition control and multi-condition joint control generation tasks, its test set includes samples covering both settings. This design ensures independence from the training set and prevents data leakage, thereby avoiding biased evaluation in conditional control tasks.
IV-B Evaluation Metrics